# itemizestditemize

### v1

@@ Abstract
We present generative predictive control (GPC), a learning control framework that (i) clones a generative diffusion-based policy from expert demonstrations, (ii) trains a predictive action-conditioned world model from both expert demonstrations and random explorations, and (iii) synthesizes an online planner that ranks and optimizes the action proposals from (i) by looking ahead into the future using the world model from (ii). Crucially, we show that conditional video diffusion allows learning (near) physics-accurate visual world models and enable robust visual foresight. Focusing on planar pushing with rich contact and collision, we show GPC dominates behavior cloning across state-based and vision-based, simulated and real-world experiments.
Figure 1: Generative Predictive Control (GPC). Given visual (or state) observations, a generative policy—trained from behavior cloning—proposes an action chunk. Conditioned on the action chunk and past observations, a predictive world model—trained from diverse trajectories—predicts the future observations in “imagination”. In (a) GPC-RANK, the policy outputs multiple action proposals and for each proposal the world model predicts its future and associated reward. The proposal with highest reward in executed. In (b) GPC-OPT, the policy outputs a single action proposal and multiple gradient ascent steps are performed to optimize the action proposal for higher reward. Both GPC-RANK and GPC-OPT dominate pure behavior cloning.
Equal contribution
I Introduction
@@
  • Predictive control. With access to a predictive dynamical model of the robot and its physical environment, optimal control (OC) and (deep) reinforcement learning (RL) are the workhorses for designing performant controllers in locomotion [57], manipulation [4], and aviation [1, 31]. While OC typically designs controllers based on analytic dynamical models and online optimization [30], RL can learn a policy via interactions with a complicated and black-box dynamical model [49], either through simulation or in the real world. The nature of predictive control is to look ahead, by anticipating how the environment will change and how much reward will be collected.
  • Generative control. Motivated by the success of generative modeling in vision and language, generative control—learning a controller from expert demonstrations using behavior cloning with generative models—is rapidly rising to prominence [15, 52]. Due to the abundance of demonstrations and the unprecedented distribution learning capacity of generative models, generative control has led to robot foundation policies capable of solving diverse tasks from vision and language [21]. The nature of generative control is to look back, making decisions based on expert policies in the past.
Each paradigm has its advantages and limitations. When dynamics modeling is accurate and tractable, predictive control delivers precision and robustness (e.g., Boston Dynamics parkour and SpaceX rocket landing). However, in many cases (e.g., manipulation with contact [35] and deformable objects [65], vision-based control [48]), the dynamics can be inaccurate or computationally expensive for optimization (as in OC) and interaction (as in RL). Generative control, on the other hand, circumvents the need for careful dynamics modeling and directly learns the policy via generative modeling of the expert demonstrations. However, generative control suffers from distribution shifts (due to its supervised learning nature [38, 21]) and lack of precision and robustness (i.e., they can do lots of tasks but do few of them as well as specialized controllers [19, 10]). Moreover, whether the “scaling law” can address these issues is questionable [7, 63, 62].
@@
Can we establish a similar robot control paradigm that both looks back (i.e., generative) and looks ahead (i.e., predictive)?
Contribution. We propose generative predictive control (GPC), a framework that combines a generative robot policy with a predictive dynamical model of the world (i.e., a world model [22]). GPC contains three modules (cf. Fig. 1):
  • Generative policy training. Given expert demonstrations, GPC uses generative modeling to train a policy that approximates the conditional distribution of expert “action chunks” given past observations.
  • Predictive world modeling. A world model simulates future observations conditioned on action chunks and past observations. One would naturally want to learn the world model using the same expert demonstrations as in policy training. However, such a world model can only simulate dynamics of the expert policy and thus is not effective for correcting policy error (see §IV). We show that adding “random exploration” data—cheaper to collect—enriches the data distribution and yield a world model that strengthens the policy at test time. For state-based world modeling, a simple MLP is sufficient to learn accurate dynamics. For vision-based world modeling, we capitalize on conditional video diffusion [25, 6] to predict action-conditioned future observations.
  • Online planning. Inspired by classical model predictive control, and assuming a reward model is available, GPC uses two lightweight online algorithms to strengthen the generative policy with the predictive world model. The first one, GPC-RANK, samples many action proposals from the policy, unrolls the future conditioned on each proposal, and selects the proposal with highest reward. The second one, GPC-OPT, treats the action proposal from the policy as a “warmstart” and optimizes it to maximize reward using gradient ascent, a technique known as single shooting [16]. These two algorithms can be easily combined to perform gradient-based optimization on multiple action proposals (warmstarts) and select the best one.
@@
Novelty. Among the three modules of GPC, the first one is not new and has become the common practice in robot policy learning [15, 52]. The novelty of GPC lies in learning a predictive world model—crucially, using random exploration data—to strengthen the generative policy with online planning (as opposed to, e.g., offline RL [23, 28, 56, 17, 44]). Moreover, different from previous works leveraging generative models for world modeling, GPC learns a world model that is explicitly action-conditioned. In other words, the world model is “controllable” by the robot’s actions. We show such a controllable world model enables robust real-world robotic manipulation.
@@
Paper organization. We formalize the overall pipeline of GPC in §II and detail the approaches for state-based and vision-based world modeling in §III. We present experimental results in §IV and conclude in §VI. Review of related works and how GPC differs from them are described in §V.
II Overview of Generative Predictive Control
@@
Let $I_{t}$ be the information vector summarizing the state of the robot and the environment up to time step $t$. For state-based robot control, $I_{t}:=x_{t}$ is the (low-dimensional) state of the robot and its environment (e.g., poses); for vision-based robot control, $I_{t}:=o_{t-H:t}$ is the history of (high-dimensional) visual observations where each $o_{t}$ is a single (or multi-view) image(s). Let $a_{t}$ be the robot action at time $t$ and $a_{t:t+T}$ be an action chunk of length $T+1$. Our goal is to design a robot control policy that decides $a_{t:t+T}$ based on $I_{t}$ to solve certain tasks (e.g., described by language instructions).
Let us formalize the three modules of GPC.
@@
Generative policy training. We ask humans to teleoperate robots to solve the tasks. The expert demonstrations are then divided into clips of data containing pairs of state information and action chunks, i.e., a dataset ${\cal D}_{\text{expert}}^{P}=\{I_{t}^{i}\leftrightarrow a_{t:t+T}^{i}\}_{i=1}^{N_{\text{expert}}}$ with number of samples $N_{\text{expert}}$. Policy learning becomes a supervised learning problem with input $I_{t}$ and output $a_{t:t+T}$. [^1] It is also common to include language instructions as input to the policy as in vision-language-action (VLA) models [32]. GPC can generalize to VLAs in a straightforward way. We exclude language here to focus on motion planning. In GPC, we follow the diffusion policy learning framework [15]. Specifically, we feed $I_{t}$ into a network that parametrizes the score function of the conditional distribution $p(a_{t:t+T}|I_{t})$, using which Gaussian noise vectors are gradually diffused to expert action chunks. The score function network is trained using denoising prediction [24]. Policy learning leads to a policy network ${\cal P}(\cdot)$ that predicts action chunks $a_{t:t+T}$ given state information $I_{t}$. It is important to note that ${\cal P}(\cdot)$ is stochastic and effectively a sampler from the (approximate) conditional distribution $p(a_{t:t+T}|I_{t})$, meaning that given $I_{t}$, we can draw many possible action chunks $(a_{t:t+T}^{(1)},\dots,a_{t:t+T}^{(K)})$ from ${\cal P}(\cdot)$. We call each action chunk $a_{t:t+T}$ an action proposal.
@@
Predictive world modeling. The fact that ${\cal P}(\cdot)$ is stochastic leads to a natural question: which one of the $K$ action proposals should be selected? The world model precisely aims to answer this question by predicting the future consequences of each action proposal. To build the world model, we need pairs of $(I_{t},a_{t:t+T})$ (current state information plus an action chunk) and $I_{t+1:t+T+1}$ (future state information). The expert demonstrations used for policy training already contain such pairs, i.e., a dataset ${\cal D}_{\text{expert}}^{W}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{expert}}}$. Nevertheless, as we show in §IV, if we train the world model only using ${\cal D}_{\text{expert}}^{W}$, the predictive power of the world model is insufficient to guide action proposal selection. One can collect more expert demonstrations to enlarge the datasets, but it can be expensive. Instead, we posit that random exploration data are not only easier to collect, but also rich in dynamics. Therefore, we collect another exploration dataset ${\cal D}^{W}_{\text{explore}}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{explore}}}$ where humans (or any other controllers) are asked to randomly perturb the system without solving any tasks. In control theory terminology, the goal of this is to perform “system identification” with “sufficient excitation” [37]. We combine the two datasets ${\cal D}^{W}:={\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}$ to train the world model. Details of the design of the world model architecture are deferred to §III. For now, the reader can assume we obtain a model ${\cal W}(\cdot)$ that predicts the future $I_{t+1:t+T+1}$ conditioned on the current state information and the action proposal $(I_{t},a_{t:t+T})$.
Online planning. We now formalize the two online planning algorithms, namely GPC-RANK and GPC-OPT, that combine the policy ${\cal P}(\cdot)$ and the world model ${\cal W}(\cdot)$ (cf. Fig. 1). We assume a reward model ${\cal R}(\cdot):I_{t+1:t+T+1}\rightarrow r_{t}$ is available that predicts the reward given the future state information, e.g., a small neural network that is separately trained.
@@
    GPC-RANK is based on the simple intuition to pick the action proposal with highest reward. Formally, we sample $K$ action proposals from the policy
@@
    and output the best proposal by passing them through the world model ${\cal W}(\cdot)$ and computing respective rewards:
@@
    where we used the notation $\pi(\cdot)$ to denote the online policy in contrast to the offline policy ${\cal P}(\cdot)$. The underlying assumption of GPC-RANK is that, due to distribution shifts, the optimal action chunk at test time may not have the highest likelihood in ${\cal P}(\cdot)$, and thus may need more samples to be discovered. The world model—assuming it is learned from a richer distribution because of random exploration—will pick up the optimal action chunk through (2). A nice property of GPC-RANK is that it can be easily parallelized.
@@
    GPC-OPT takes a different perspective by directly solving the reward maximization problem given the world model:
@@
    where the action chunk are viewed as decision variables. The difficulty is, however, problem (3) can be highly nonconvex due to the nonlinearity and nonsmoothness of ${\cal W}(\cdot)$ and ${\cal R}(\cdot)$. Fortunately, the pretrained policy ${\cal P}(\cdot)$ comes to rescue because it narrows down the search space by providing good initializations to solve (3), i.e., warmstarts. Formally, we sample an action chunk $\hat{a}_{t:t+T}$ from ${\cal P}(\cdot)$ and perform $M$ steps of gradient ascent starting from $\hat{a}_{t:t+T}$ (assuming ${\cal R}(\cdot)$ is differentiable):
    $\hat{a}_{t:t+T}^{(0)}:=\hat{a}_{t:t+T}\sim{\cal P}(I_{t}),\text{ and for }\ell=1,\dots,M\text{ do:}$:
    $\hat{a}_{t:t+T}^{(\ell)}=\hat{a}_{t:t+T}^{(\ell-1)}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)})),$ (4)
    where $\eta^{(\ell)}>0$ are step sizes. The online policy outputs the final optimized action chunk
@@
    Since the gradients in ((ii)) are readily available from automatic differentiation, gradient ascent can be carried out using off-the-shelf optimizers like ADAM. Readers familiar with the optimal control literature will recognize GPC-OPT is the classical single shooting algorithm [16]. Compared to GPC-RANK which evaluates the world model $K$ times, GPC-OPT evaluates the world model $M$ times. The former can be parallelized but the latter cannot.

### v2

@@ Abstract
We present generative predictive control (GPC), a learning control framework that (i) clones a generative diffusion-based policy from expert demonstrations, (ii) trains a predictive action-conditioned world model from both expert demonstrations and random explorations, and (iii) synthesizes an online planner that ranks and optimizes the action proposals from (i) by looking ahead into the future using the world model from (ii). Across a variety of robotic manipulation tasks, we demonstrate that GPC consistently outperforms behavior cloning in both state-based and vision-based settings, in simulation and in the real world.
Figure 1: Generative Predictive Control.
1 Introduction
@@
Given access to a predictive model of the robot and its environment, predictive control, e.g., optimal control (OC) and reinforcement learning (RL), serves as the primary tool for designing high-performance controllers in locomotion [1], manipulation [2], and aviation [3, 4]. OC designs controllers using analytic models and online optimization [5], while RL learns policies through interaction with complex, often black-box dynamics [6], either in simulation or the real world. At their core, predictive control methods look ahead, anticipating environmental changes and future rewards.
Inspired by the success of generative modeling in vision and language, generative control—learning controllers from expert demonstrations using behavior cloning with generative models—is rapidly gaining prominence [7, 8]. Leveraging abundant demonstrations and the powerful distribution learning capabilities of generative models, generative control has enabled the creation of robot foundation policies that can solve a wide range of tasks from vision and language inputs [9, 10]. At its core, generative control looks back, basing decisions on past expert behaviors.
Each paradigm has its advantages and limitations. When dynamics and reward modeling is accurate and tractable, predictive control delivers precision and robustness (e.g., Boston Dynamics parkour and SpaceX rocket landing). However, in many cases (e.g., manipulation with contact [11] and deformable objects [12], vision-based control [13]), the dynamics can be inaccurate or computationally expensive for optimization (as in OC) and interaction (as in RL). Generative control, on the other hand, circumvents the need for careful dynamics and reward modeling and directly learns the policy via generative modeling of the expert demonstrations. However, generative control suffers from distribution shifts (due to its supervised learning nature [14, 9]) and lack of precision and robustness (i.e., they can do lots of tasks but do few of them as well as specialized controllers [15, 16]). Moreover, whether the “scaling law” can address these issues remains unclear [17, 18, 19].
@@
Can we establish a similar robot control paradigm that both looks back and looks ahead?
Contribution. We propose generative predictive control (GPC), a framework that combines a generative policy with a predictive world model [20]. GPC consists of three modules (cf. Fig. 1):
  • Generative policy training. Given expert demonstrations, GPC uses generative modeling, particularly diffusion models, to train a policy that approximates the conditional distribution of expert “action chunks” given past observations.
  • Predictive world modeling. A world model simulates future observations conditioned on action chunks and past observations. One would naturally want to learn the world model using the same expert demonstrations as in policy training. However, such a world model can only simulate dynamics of the expert policy and thus is not effective for correcting policy error (see §4). We show that adding “random exploration” data—cheaper to collect—enriches the data distribution and yields a world model that strengthens the policy at test time. For state-based world modeling, a simple MLP is sufficient to learn accurate dynamics. For vision-based world modeling, we capitalize on conditional video diffusion [21, 22] to predict action-conditioned future observations.
  • Online planning. Inspired by classical model predictive control, and assuming a reward model is available, GPC uses two lightweight online algorithms to strengthen the generative policy with the predictive world model. The first one, GPC-RANK, samples many action proposals from the policy, unrolls the future conditioned on each proposal, and selects the proposal with highest reward. The second one, GPC-OPT, treats the action proposal from the policy as a “warmstart” and optimizes it to maximize reward using gradient ascent, in the spirit of single shooting [23]. These two algorithms can be combined to perform gradient-based optimization on multiple action proposals and select the best one. Moreover, we show that the reward model can either be learned from expert demonstrations or implemented using a zero-shot VLM, making GPC broadly applicable across diverse tasks.
@@
Paper organization. We formalize the overall pipeline of GPC in §2 and detail the approaches for state-based and vision-based world modeling in §3 and Appendix C. We present experimental results in §4 and conclude in §5. More results can be found in Appendix F and https://computationalrobotics.seas.harvard.edu/GPC/. Related works are reviewed in Appendix G.
2 Overview of Generative Predictive Control
@@
Let $I_{t}$ be the information vector summarizing the state of the robot and the environment up to time step $t$. For state-based robot control, $I_{t}:=x_{t}$ is the (low-dimensional) state of the robot and its environment (e.g., poses); for vision-based robot control, $I_{t}:=o_{t-H:t}$ is the history of (high-dimensional) visual observations where each $o_{t}$ is a single (or multi-view) image(s). Let $a_{t}$ be the robot action at time $t$ and $a_{t:t+T}$ be an action chunk of length $T+1$. Our goal is to design a robot control policy that decides $a_{t:t+T}$ based on $I_{t}$ to solve certain tasks (e.g., described by language instructions). Let us formalize the three modules of GPC.
@@
Generative policy training. We ask humans to teleoperate robots to solve tasks, generating expert demonstrations that are segmented into clips of state-action pairs, forming a dataset ${\cal D}_{\text{expert}}^{P}=\{I_{t}^{i}\leftrightarrow a_{t:t+T}^{i}\}_{i=1}^{N_{\text{expert}}}$. Policy learning then reduces to supervised learning with input $I_{t}$ and output $a_{t:t+T}$. In GPC, we adopt the diffusion policy learning framework [7], where $I_{t}$ is fed into a network that parametrizes the score function of $p(a_{t:t+T}|I_{t})$. Gaussian noise vectors are gradually denoised into expert action chunks following [24]. This yields a stochastic policy network ${\cal P}(\cdot)$ that samples action chunks $a_{t:t+T}$ given $I_{t}$. We refer to each sampled action chunk as an action proposal.
@@
Predictive world modeling. The stochastic nature of ${\cal P}(\cdot)$ raises a key question: which of the many action proposals should be selected? The world model addresses this by predicting the future outcomes of each proposal. To train it, we use pairs of $(I_{t},a_{t:t+T})$ and $I_{t+1:t+T+1}$ from the expert dataset ${\cal D}_{\text{expert}}^{W}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{expert}}}$, already available from policy training. However, as shown in §4, training solely on ${\cal D}_{\text{expert}}^{W}$ leads to limited predictive ability. While more expert data could help, it can be costly. Instead, we collect an additional exploration dataset ${\cal D}^{W}_{\text{explore}}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{explore}}}$, where humans (or other controllers) randomly perturb the system without solving tasks. This approach, inspired from “system identification” with “sufficient excitation” in control theory [25], enriches the dynamics. We combine both datasets as ${\cal D}^{W}:={\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}$ to train the world model. Architectural details are in §3; for now, we assume access to a model ${\cal W}(\cdot)$ that predicts $I_{t+1:t+T+1}$ from $(I_{t},a_{t:t+T})$.
Online planning. We now formalize the two online planning algorithms, namely GPC-RANK and GPC-OPT, that combine the policy ${\cal P}(\cdot)$ and the world model ${\cal W}(\cdot)$. We assume a reward model ${\cal R}(\cdot):I_{t+1:t+T+1}\rightarrow r_{t}$ is available that predicts the reward given the future state information, e.g., a small neural network that is separately trained or foundation models (i.e., vision-language models (VLMs)) capable of zero-shot selection of the most promising future state. A trained reward predictor is appropriate in scenarios where the reward can be explicitly defined—particularly in numerical terms—and it is differentiable, enabling gradient-based action optimization. However, some tasks involve rewards that are difficult or even infeasible to specify. In such cases, VLMs can serve as surrogate reward predictors by directly selecting the most suitable action proposal, given the predicted outcomes and task description as input. This approach introduces greater flexibility to our GPC method, extending its applicability to a broader range of tasks.
@@
    GPC-RANK is based on the intuition to pick the action proposal with highest reward. Formally, we sample $K$ action proposals from the policy
@@
    pass them through the world model ${\cal W}(\cdot)$, and select the one with highest reward:
@@
    where we used the notation $\pi(\cdot)$ to denote the online policy in contrast to the offline policy ${\cal P}(\cdot)$. Due to distribution shifts, the optimal action chunk at test time may not correspond to the highest-likelihood sample from ${\cal P}(\cdot)$ and thus may require more samples to uncover. By training on a richer distribution via random exploration, the world model helps select the optimal chunk through (2). A key advantage of GPC-RANK is its easy parallelizability.
@@
    GPC-OPT directly solves the reward maximization problem given the world model:
@@
    treating the action chunk as decision variables. However, (3) is often highly nonconvex due to the complexity of ${\cal W}(\cdot)$ and ${\cal R}(\cdot)$. To address this, we leverage the pretrained ${\cal P}(\cdot)$ to warm-start optimization. We sample $\hat{a}_{t:t+T}$ from ${\cal P}(I_{t})$ and perform:
    $\hat{a}_{t:t+T}^{(0)}:=\hat{a}_{t:t+T}\sim{\cal P}(I_{t}),\quad\hat{a}_{t:t+T}^{(\ell)}=\hat{a}_{t:t+T}^{(\ell-1)}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)})),$ (4)
    for $\ell=1,\dots,M$, with step sizes $\eta^{(\ell)}>0$. The final optimized action chunk is:
@@
    Gradients are obtained via automatic differentiation, and updates can be performed with optimizers like ADAM. GPC-OPT resembles the classical single shooting method [23]. Unlike GPC-RANK, which evaluates the world model $K$ times in parallel, GPC-OPT evaluates it $M$ times sequentially.